A wall that is 12 feet wide and 8 feet high must be painted. There are 2 windows measuring 3 by 5 feet on the wall. There is also a door that is 4 by 7 feet. What percent of the wall can be painted?

Answer by checkley75(3666)

12*8=96 ft^2 is the total wall area.
2*3*5=30
4*7=28
96-30-28=38 ft^2 to be painted.
38/96=39.6 % needs to be painted.